# Exploring the Design Space of Diffusion Bridge Models via Stochasticity Control

### v1

@@ Appendix A Proofs
Theorem 3.
Suppose the transition kernel of a diffusion process is given by $p_{t|0,T}(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$, then the evolution of conditional probability $q(\mathbf{X}_{t}|\mathbf{x}_{T})$ has a class of time reverse sampling SDEs of the form:
$d\mathbf{X}_{t}$ $=\left[\dot{\alpha}_{t}\hat{\mathbf{x}}_{0}+\dot{\beta}_{t}\mathbf{x}_{T}-(\dot{\gamma}_{t}\gamma_{t}+\epsilon_{t})\nabla_{\mathbf{X}_{t}}\log p_{t}(\mathbf{X}_{t}|\mathbf{x}_{T})\right]dt+\sqrt{2\epsilon_{t}}d\mathbf{W}_{t}\quad\mathbf{X}_{T}=\mathbf{x}_{T}.$ (34)
Proof.
Recall Eqs. (10) 11 and Lemma 2,
Next we take the reparameterized score 15 into 35:
Theorem 4.
Let $(\mathbf{x}_{0},\mathbf{x}_{T})\sim\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$, $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}|\mathbf{x}_{0},\mathbf{x}_{T})$, Given the transition kernel: $p(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I}\right)$, if $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ is a denoiser function that minimizes the expected $L_{2}$ denoising error for samples drawn from $\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$:
@@
Appendix B Reframing previous methods in our framework
@@ B.1 DDBM-VE
Compare Eqs. (56) and 57 with Lemma 1. We only need to prove:
@@ B.3 EDM
Recall 11, the ODE formulation is given by:
@@
Denoising score matching. The score remarameterization in EDM is the same as ours in Eq. 15. Let $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$, then the score reparameterization in Eq. 15 is given by:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{X}_{t})\approx\frac{\hat{\mathbf{x}}_{0}-\mathbf{x}_{t}}{\sigma_{t}^{2}}.$ (85)
Sampling SDEs with stochasticity added. Recall Theorem 3, as $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$, then the SDE has the form:
@@ B.4 I2SB
Using discretization 20:
@@
Appendix C Experiment Details
@@
Architecture. We maintain the architecture and parameter settings consistent with (Linqi Zhou et al., 2023), utilizing the ADM model (Dhariwal & Nichol, 2021) for $64\times 64$ resolution, modifying the channel dimensions from $192$ to $256$ and reducing the number of residual blocks from three to two. Apart from these changes, all other settings remain identical to those used for $64\times 64$ resolution.
@@
Training. We include additional pre- and post-processing steps: scaling functions and loss weighting, the same ingredient as (Karras et al., 2022). Let $D_{\theta}(\mathbf{x}_{t},\mathbf{x}_{T},t)=c_{\mathrm{skip}}(t)\mathbf{x}_{t}+c_{\mathrm{out}(t)}(t)F_{\theta}(c_{\mathrm{in}}(t)\mathbf{x}_{t},c_{\mathrm{noise}}(t))$, where $F_{\theta}$ is a neural network with parameter $\theta$, the effective training target with respect to the raw network $F_{\theta}$ is: $\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{0},\mathbf{x}_{T},t}\left[\lambda\|c_{\mathrm{skip}}(\mathbf{x}_{t}+c_{\mathrm{out}}{F_{\theta}}(c_{\mathrm{in}}\mathbf{x}_{t},c_{\mathrm{noise}})-\mathbf{x}_{0}\|^{2}\right]$. Scaling scheme are chosen by requiring network inputs and training targets to have unit variance $(c_{\mathrm{in}},c_{\mathrm{out}})$, and amplifying errors in $F_{\theta}$ as little as possible. Following reasoning in (Linqi Zhou et al., 2023),
@@
$c_{\mathrm{in}}(t)=\frac{1}{\sqrt{\alpha_{t}^{2}\sigma_{0}^{2}+\beta_{t}^{2}\sigma_{T}^{2}+2\alpha_{t}\beta_{t}\sigma_{01}+\gamma_{t}^{2}}},\quad c_{\mathrm{skip}}(t)=(\alpha_{t}\sigma_{0}^{2}+\beta_{t}\sigma_{01})*c_{\mathrm{in}}^{2},$ (93)
@@
$c_{\mathrm{out}}(t)=\sqrt{\beta_{t}^{2}\sigma_{0}^{2}\sigma_{1}^{2}-\beta_{t}^{2}\sigma_{01}^{2}+\gamma_{t}^{2}\sigma_{0}^{2}}c_{\mathrm{in}},\quad\lambda=\frac{1}{c_{\mathrm{out}}^{2}},\quad c_{\mathrm{noise}}(t)=\frac{1}{4}\log{(t)},$ (94)
@@
We set $\sigma_{0}=\sigma_{T}=0.5,\sigma_{0T}=\sigma_{0}^{2}/2$ for all training sessions. Other setting are shown in Table 5.
Table 5: Training setting
@@
$\eta$ 00 $0.5$
$\gamma_{\max}$ $0.125$ $0.125$
Setting GPU 1 A6000 48G 1 H100 96G
Batch size 32 200
Learning rate $1\times 10^{-5}$ $1\times 10^{-4}$
Iterations $9\times 10^{6}$ $1\times 10^{6}$
Training time 42 days 11 days
Sampling. We use the same timesteps distributed according to EDM (Karras et al., 2022): $(t_{\max}^{1/\rho}+\frac{i}{N}(t_{\min}^{1/\rho}-t_{\max}^{1/\rho}))^{\rho}$, where $t_{\min}=0.01$ and $t_{\max}=1-10^{-4}$. The best performance achieved by setting $\rho=0.6$.
Licenses
@@
  • DIODE-Outdoor Vasiljevic et al. (2019): MIT license.
Appendix D Additional visualizations
@@
Figure 6: DDBM model and SDB sampler ($\eta=0.3$, NFE=$20$, FID=$4.12$). Samples for DIODE dataset (conditoned on depth images).
Figure 7: SDB model and sampler ( $\gamma_{\max}=0.125$, $\eta=1$, $b=0$, NFE=$5$, FID=$0.89$).
@@
Figure 9: SDB model and sampler ($\gamma_{\max}=0.125$, $b=0.5$, NFE=40, FID = 1.71, AFD=9.52).

### v2

@@ Appendix B Proofs
Lemma B.3.
Let $(\mathbf{x}_{0},\mathbf{x}_{T})\sim\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$, $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}|\mathbf{x}_{0},\mathbf{x}_{T})$, Given the transition kernel: $p(\mathbf{x}_{t}\mid\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}\left(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbb{I}\right)$, if $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},\mathbf{x}_{T},t)$ is a denoiser function that minimizes the expected $L_{2}$ denoising error for samples drawn from $\pi_{0}(\mathbf{x}_{0},\mathbf{x}_{T})$:
@@
Proof of Prop. 4.1.
Proof.
Recall Eqs. (21) (22) and Lemma B.2,
Next we take the reparameterized score in Eq. 31 into Eq. 45:
Appendix C Reframing previous methods in our framework
Table 8: Specify design choices for different model families. In the implementation, $\sigma_{t}=t$ for EDM, $\sigma_{t}=t,a_{t}=1$ for DDBM-VE, $\sigma_{t}=\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{\min}t}-1}$ and $a_{t}=1/\sqrt{e^{\frac{1}{2}\beta_{d}t^{2}+\beta_{\min}t}}$ for DDBM-VP, where $\beta_{d}$ and $\beta_{\min}$ are parameters. We include details and proofs in Appendix C.
I2SB DDBM DBIM EDM Ours
Transition kernel $\alpha_{t}$ $1-\sigma_{t}^{2}/\sigma_{T}^{2}$ $a_{t}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $a_{t}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $1$ $1-t$
$\beta_{t}$ $\sigma_{t}^{2}/\sigma_{T}^{2}$ ${a_{T}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t})$ ${a_{T}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t})$ 00 $t$
$\gamma_{t}^{2}$ $\sigma_{t}^{2}(1-\sigma_{t}^{2}/\sigma_{T}^{2})$ $\sigma_{t}^{2}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $\sigma_{t}^{2}(1{-}{a_{T}^{2}\sigma_{t}^{2}}/(\sigma_{t}^{2}a_{t}^{2}))$ $\sigma_{t}^{2}$ $\frac{\gamma_{\max}^{2}}{4}{t(1-t)}$
Sampling SDEs $\epsilon_{t}$ $\frac{\gamma_{t-\Delta t}^{2}\beta_{t}^{2}-\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{2\beta_{t}^{2}\Delta t}$ $\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$$\begin{cases}\frac{\gamma_{t-\Delta t}^{2}}{2\Delta t},&t=0\\ 0,&t\neq 0\end{cases}$ $\bar{\beta_{t}}\sigma_{t}^{2}$ $\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$
$\eta=0$ or $\eta=1$ $\eta\in[0,1]$
Base distribution $\pi_{T}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}$ $\pi_{\mathrm{cond}}*\mathcal{N}(0,b^{2}\mathbb{I})$
Discretization - Euler Euler Euler Heun Euler
Eq. (12) Eq. (13) Eq. (12) - Eqs. (13) and (12)
@@ C.1 DDBM-VE
Compare Eqs. (51) and 52 with Lemma B.1. We only need to prove:
@@ C.3 EDM
Recall 22, the ODE formulation is given by:
@@
Sampling SDEs with noise added. Recall Proposition 4.1, as $\alpha_{t}=1,\beta_{t}=0,\gamma_{t}=\sigma_{t}$, then the SDE has the form:
@@ C.4 I2SB
${x}_{t-\Delta t}$ $=(\alpha_{t-\Delta t}-\alpha_{t}\frac{\beta_{t-\Delta t}}{\beta_{t}})\hat{{x}}_{0}+$
$\frac{\beta_{t-\Delta t}}{\beta_{t}}{x}_{t}+\sqrt{\gamma_{t-\Delta t}^{2}-\frac{\beta_{t-\Delta t}^{2}\gamma_{t}^{2}}{\beta_{t}^{2}}}\bar{{z}}_{t}$ (82)
Using discretization in Sec. C.4:
@@
Appendix D Additional design guideline
Figure 7: Ablation study on the shape of $\gamma_{t}$.
$\alpha_{t}$ and $\beta_{t}$. Theoretically, $\alpha_{t}$ and $\beta_{t}$ can be freely designed, and future work may explore alternative design choices. However, in this paper, we focus on the simple case where $\alpha_{t}=1-t$ and $\beta_{t}=t$. The rationale is as follows: consider the scenario where $\alpha_{t}=1-\beta_{t}$, which represents an interpolation along the line segment between $x_{0}$ and $x_{1}$. For the path $p_{t}^{(1)}(x)=\mathcal{N}((1-\beta_{t})x_{0}+\beta_{t}x_{1},\gamma_{t}^{2}\mathbb{I})$, where $\beta_{t}$ is invertible, it is straightforward to construct another path $p_{t}^{(2)}(x)=\mathcal{N}((1-t)x_{0}+tx_{1},\gamma_{\beta_{t}^{-1}}^{2}\mathbb{I})$, which achieves the same objective function but uses a different distribution of $t$ during training. Based on this equivalence, setting $\alpha_{t}=1-t$ and $\beta_{t}=t$ is a reasonable choice.
Figure 8: Sampling paths with dfferent choices of $\gamma_{t}$. As $\gamma_{t}$ extreamly low, e.g, $\gamma_{\max}=0.025$, the model will be failed to construct details of images.
The shape of $\gamma_{t}$. We conducted an ablation study on $\gamma_{t}$ with different shapes. Specifically, we assumed $\gamma_{t}$ has the form $\gamma_{t}=2\gamma_{\max}\sqrt{t^{k}(1-t^{k})}$, as shown in Fig. 7, $\gamma_{t}$ will have different shape as we set different $k$. The results indicate that the best performance is achieved when $k=1$, which is the exact setting used in this paper.
$\gamma_{\max}$. Our ablation studies on $\gamma_{\max}$ demonstrate that the optimal values of $\gamma_{\max}$ are approximately $0.125$ or $0.25$. Furthermore, the sampling paths corresponding to different choices of $\gamma_{t}$ are shown in Fig. 8. Adding an appropriate amount of noise to the transition kernel helps in constructing finer details.
Figure 9: An illustration of design choices of transition kernels and how they affect the I2I translation process. $\alpha_{t}$ and $\beta_{t}$ define the interpolation between two images, while $\gamma_{t}$ controls the noise added to the process. ntuitively, the DDBM-VE model introduces excessive noise in the middle stages, which is unnecessary for effective image translation and may explain its poor performance. In contrast, our Linear path results in a symmetrical noise schedule, ensuring a more balanced process. On the other hand, the DDBM-VP path adds more noise near $\mathbf{x}_{T}$, , indicating that during training, more computational resources are focused around $\mathbf{x}_{0}$.
Figure 10: Sampling path with dfferent choices of $\epsilon_{t}$. As $\epsilon_{t}=0$, the generated images lack details, as $\epsilon_{t}$ too large, the sampled images are over-sharpening. The best choices of $\epsilon_{t}$ are around $\epsilon_{t}=0.8$ and $\epsilon_{t}=1.0$.
$\epsilon_{t}$. We use the setting $\epsilon_{t}=\eta\left(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2}\right).$ The ablation studies on $\epsilon_{t}$ demonstrate that the optimal choice of $\eta$ for the DDBM-VP model is approximately $0.3$, while the best choice for the ECSI model with a Linear Path is around $1.0$. Additionally, we present sample paths and generated images under different $\eta$ settings to illustrate heuristic parameter tuning techniques. The results are shown in Figures 10, 11, and 12. Too small a value of $\eta$ results in the loss of high-frequency information, while too large a value of $\eta$ produces over-sharpened and potentially noisy sampled images.
Figure 11: Comparison of sampled images with different $\epsilon_{t}$ for ECSI model, where $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$, $\gamma_{\max}=0.25$, $b=0$.
Figure 12: Comparison of sampled images with different $\epsilon_{t}$ for DDBM-VP pretrained model, where $\epsilon_{t}=\eta(\gamma_{t}\dot{\gamma}_{t}-\frac{\dot{\alpha}_{t}}{\alpha_{t}}\gamma_{t}^{2})$.
Appendix E Impact Statement
Our method can improve image translation and solving inverse problem, which may benefit applications in medical imaging. However, it is important to note that as with many generative and restoration models, our method could be misused for malicious image manipulation.
Appendix F Experiment Details
@@
Architecture. We maintain the architecture and parameter settings consistent with [17], utilizing the ADM model [10] for $64\times 64$ resolution, modifying the channel dimensions from $192$ to $256$ and reducing the number of residual blocks from three to two. Apart from these changes, all other settings remain identical to those used for $64\times 64$ resolution.
@@
Training. We include additional pre- and post-processing steps: scaling functions and loss weighting, the same ingredient as [16]. Let $D_{\theta}(\mathbf{x}_{t},\mathbf{x}_{T},t)=c_{\mathrm{skip}}(t)\mathbf{x}_{t}+c_{\mathrm{out}(t)}(t)F_{\theta}(c_{\mathrm{in}}(t)\mathbf{x}_{t},c_{\mathrm{noise}}(t))$, where $F_{\theta}$ is a neural network with parameter $\theta$, the effective training target with respect to the raw network $F_{\theta}$ is: $\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{0},\mathbf{x}_{T},t}\left[\lambda\|c_{\mathrm{skip}}(\mathbf{x}_{t}+c_{\mathrm{out}}{F_{\theta}}(c_{\mathrm{in}}\mathbf{x}_{t},c_{\mathrm{noise}})-\mathbf{x}_{0}\|^{2}\right]$. Scaling scheme are chosen by requiring network inputs and training targets to have unit variance $(c_{\mathrm{in}},c_{\mathrm{out}})$, and amplifying errors in $F_{\theta}$ as little as possible. Following reasoning in [17],
@@
$c_{\mathrm{in}}(t)=\frac{1}{\sqrt{\alpha_{t}^{2}\sigma_{0}^{2}+\beta_{t}^{2}\sigma_{T}^{2}+2\alpha_{t}\beta_{t}\sigma_{0T}+\gamma_{t}^{2}}},\quad c_{\mathrm{skip}}(t)=(\alpha_{t}\sigma_{0}^{2}+\beta_{t}\sigma_{0T})*c_{\mathrm{in}}^{2},$ (88)
@@
$c_{\mathrm{out}}(t)=\sqrt{\beta_{t}^{2}\sigma_{0}^{2}\sigma_{1}^{2}-\beta_{t}^{2}\sigma_{0T}^{2}+\gamma_{t}^{2}\sigma_{0}^{2}}c_{\mathrm{in}},\quad\lambda=\frac{1}{c_{\mathrm{out}}^{2}},\quad c_{\mathrm{noise}}(t)=\frac{1}{4}\log{(t)},$ (89)
@@
We note that TrigFlow [22], adopts the same score reparameterization and pre-conditioning techniques. It can be considered a special case of our framework by setting $\alpha_{t}=\cos(t)$, $\beta_{t}=0$, $\gamma_{t}=\sigma_{0}\sin(t)$, $t\in[0,\frac{\pi}{2}]$. In this case, $\sigma_{T}=0$, $\sigma_{0T}=0$,
$c_{\mathrm{in}}(t)$ $=\frac{1}{\sqrt{\alpha_{t}^{2}\sigma_{0}^{2}+\gamma_{t}^{2}}}=\frac{1}{\sqrt{\sin^{2}(t)\sigma_{0}^{2}+\cos^{2}(t)\sigma_{0}^{2}}}=\frac{1}{\sigma_{0}},$ (90)
$c_{\mathrm{skip}}(t)$ $=(\alpha_{t}\sigma_{0}^{2})c_{in}^{2}=\cos(t)\cdot\sigma_{0}^{2}\cdot\frac{1}{\sigma_{0}^{2}}=\cos(t),$ (91)
$c_{out}(t)$ $=\sqrt{\gamma_{t}^{2}\sigma_{0}^{2}}\cdot c_{in}=\sin(t)\sigma_{0},$ (92)
$D_{\theta}(x_{t},t)$ $=c_{\mathrm{skip}}x_{t}+c_{\mathrm{out}}F_{\theta}(c_{\mathrm{in}}x_{t},c_{\mathrm{noise}})=\cos(t)x_{t}+\sin(t)\sigma_{0}F_{\theta}(\frac{1}{\sigma_{0}},c_{\mathrm{noise}}).$ (93)
Then we recover TrigFlow.
In our implementation, we set $\sigma_{0}=\sigma_{T}=0.5,\sigma_{0T}=\sigma_{0}^{2}/2$ for all training sessions. Other setting are shown in Table 9.
Table 9: Training settings
@@
$\eta$ 00 00 $0.5$
$\gamma_{\max}$ $0.125$ $0.25$ $0.125$
Setting GPU 1 A6000 48G 1 H100 96G 1 H100 96G
Batch size 32 128 200
Learning rate $1\times 10^{-5}$ $5\times 10^{-5}$ $1\times 10^{-4}$
epochs $2078$ $2106$ $1443$
Training time 42 days 8 days 11 days
Model Dataset DIODE ($256\times 256$) DOIDE ($256\times 256$)
$\eta$ 00 00
$\gamma_{\max}$ $0.125$ $0.25$
Setting GPU 1 H100 96G 1 H100 96G
Batch size 16 16
Learning rate $2\times 10^{-5}$ $2\times 10^{-5}$
epochs $2617$ $1745$
Training time 17 days 25 days
Sampling. We use the same timesteps distributed according to EDM [16]: $(t_{\max}^{1/\rho}+\frac{i}{N}(t_{\min}^{1/\rho}-t_{\max}^{1/\rho}))^{\rho}$, where $t_{\min}=0.001$ and $t_{\max}=1-10^{-4}$. The best performance achieved by setting $\rho=0.6$ for Edges2handbags and $\rho=0.8$ for DIODE datasets.
Appendix G Licenses
@@
  • DIODE-Outdoor [36]: MIT license.
Figure 13: ECSI model and sampler ( $\gamma_{\max}=0.125$, $\eta=1$, $b=0$, NFE=$5$, FID=$0.89$).
Appendix H Additional visualizations
@@
Figure 15: DDBM model and ECSI sampler ($\eta=0.3$, NFE=$20$, FID=$4.12$). Samples for DIODE dataset (conditoned on depth images).
Figure 16: ECSI model and sampler ($\gamma_{\max}=0.25$, $\eta=1.0$, $b=0$, NFE=5, FID = 4.16).
Figure 17: ECSI model and sampler ($\gamma_{\max}=0.25$, $\eta=1.0$, $b=0$, NFE=20, FID = 3.27).
Figure 18: DDBM model and DBIM sampler (NFE=10, FID = 2.46, AFD=5.20).
@@
Figure 20: ECSI model and sampler ($\gamma_{\max}=0.125$, $b=1.0$, NFE=10, FID = 2.07, AFD=9.35).
Figure 21: DDBM model and ECSI sampler on $446$ test images. (NFE=20, FID = 52.01, AFD=5.60).
Figure 22: ECSI model and sampler on $446$ test images. ($\gamma_{\max}=0.125$, $b=0.5$, NFE=20, FID = 55.93, AFD=7.39).